OpenStudy (anonymous):

Evaluate: -8/9 + 4/9 - 2/9

jimthompson5910 (jim_thompson5910):

The denominators are the same, so you can combine the numerators over the common denominator

OpenStudy (anonymous):

Okay? how?

jimthompson5910 (jim_thompson5910):

example: 1/8 + 2/8 = (1+2)/8 1/8 + 2/8 = 3/8

jimthompson5910 (jim_thompson5910):

Notice how I combined the two numerators 1 and 2 to get 3. I placed all this over the common denominator of 8.

OpenStudy (anonymous):

Okay than do i subtract 8 from the other numbers?

jimthompson5910 (jim_thompson5910):

yes because the -8 means "subtract 8"

OpenStudy (anonymous):

so is the answer -2/9?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

ooookkkkkkaaaayyyy

OpenStudy (anonymous):

But is - 2/9 from 4/9 and -8/9

jimthompson5910 (jim_thompson5910):

-8/9 + 4/9 - 2/9 (-8 + 4 - 2)/9 I'll let you do the last step

OpenStudy (anonymous):

-6/9? If not I feel stupid..

jimthompson5910 (jim_thompson5910):

-8 + 4 = ???

jimthompson5910 (jim_thompson5910):

oh wait, nvm, you got it the next step is -6/9 and that reduces to -2/3

OpenStudy (anonymous):

-4!!!

OpenStudy (anonymous):

Oh

jimthompson5910 (jim_thompson5910):

correct, so we have -8/9 + 4/9 - 2/9 (-8 + 4 - 2)/9 (-4 - 2)/9 -6/9 -2/3

OpenStudy (anonymous):

awesome! thx

jimthompson5910 (jim_thompson5910):

sure thing